# Exploring the effect of different cosmologies on the Epoch of Reionization 21-cm signal with Polar

### v1

@@ 2.2.1 Semi-Analytic modeling of galaxies
We use L-Galaxies as described by Barrera et al. (2023), as a post-processing module of Gadget-4. This is an updated version of the publicly available L-Galaxies 2020 that was used by M23. L-Galaxies implements most major physical processes of gas cooling, star formation, galaxy mergers, supernovae feedback, black hole growth, AGN feedback, and dust attenuation. While Barrera et al. (2023) largely builds on Henriques et al. (2015), we also consider the parameters used by Henriques et al. (2020) before adapting them for our purposes. In particular, we focus on those that control star formation efficiency ($\alpha_{\rm SF}$), star formation efficiency during galaxy mergers ($\alpha_{\rm SF,burst}$), AGN accretion rate ($k_{\rm AGN}$), reheating of cold gas by star formation ($\epsilon_{\rm reheat}$, $V_{\rm reheat}$), and the energy released by each supernova ($E_{\rm SN}$).
@@ 2.2.2 Radiative transfer and the 21-cm signal
More specifically, Grizzly requires as input the gridded matter density field and dark matter halo masses from the $N$-body simulation, as well as the corresponding galactic stellar masses and stellar ages obtained with L-Galaxies. Our reference simulation has a grid of 2563 cells, resulting in a cell size of $\approx 600~{}h^{-1}~{}\rm ckpc$, but we use a range of different sizes (from 643 to 5123 cells) to assure convergence of output values in the range $512>z>$ 5. Next, we assume that the gas density scales by a factor of $\Omega_{\rm b}/\Omega_{\rm dm}$ with the dark matter density. Lastly, as done in M23, we use the Binary Population and Spectral Synthesis (BPASS; Stanway & Eldridge, 2018) code to model the Spectral Energy Distributions (SEDs) of stellar sources. In future work, we will also explore the impact of other source types such as X-ray binaries, shock-heated interstellar medium, and accreting black holes (as done in Eide et al., 2018, 2020; Ma et al., 2021).
@@
where $x_{\rm HI}$ is the fraction of neutral hydrogen, $\delta_{\rm B}$ is the fractional overdensity of baryons, $T_{\rm S}$ is the hydrogen spin temperature, $T_{\rm CMB}$ is the temperature of the CMB photons at redshift $z$, and $\Omega_{\rm m}$ is the total matter density. As done by M23, we assume $T_{\rm S}\gg T_{\rm CMB}$ which is valid when the IGM has been sufficiently heated by X-ray sources and expected to be the case in the range of redshift of interest here. Additionally, we also ignore the impact of redshift space distortions. Thus Equation 1 is reduced to
@@
We define the power spectrum of $\delta T_{\rm b}$ as:
@@
where $\delta_{D}$ is the Dirac delta function, $\delta T_{\rm b}(\mathbf{k})$ is the differential brightness temperature in Fourier space, and delimited-⟨$\langle$..$\rangle$ … ⟩ is the ensemble average. In the following, we will report our results in terms of the normalized form of the power spectrum, given by (see Peacock, 1999, for details):
@@
Table 2: List of astrophysical parameters in L-Galaxies (see Henriques et al., 2015, 2020, for more details) that can be tuned to control galaxy formation and evolution. We show the names of the parameters in column 1, and Henriques et al. (2015) and Henriques et al. (2020) values tuned to $z<3$ observations in columns 2 and 3. Our cosmological models (fiducial, $h$ high, $\sigma_{8}$ low and $\sigma_{8}$ high) tuned to UVLF observations at $z=10$ and $9$ are shown in columns 4, 5, 6 and 7 respectively. We highlight the parameters that differ from Henriques et al. (2015, 2020) values in bold. For the unconstrained case (see text for more details), we use the fiducial model parameters for all our $N$-body simulations.
@@
$Z_{\rm yield}$ 0.030 0.030 0.030 0.030 0.030 0.030
Figure 2: A and B-series averaged UVLFs for the unconstrained (top) and constrained (bottom) cases at $z=10$ and 9 for the fiducial (cyan), $h$ high (red), $\sigma_{8}$ low (purple) and $\sigma_{8}$ high (green). We also show the dust attenuated (solid) and unattenuated (dashed) UVLFs, and JWST and HST observations (Finkelstein et al., 2015; Bouwens et al., 2015, 2021; Harikane et al., 2022; Bouwens et al., 2023a; Bouwens et al., 2023b; Harikane et al., 2023; Leung et al., 2023; McLeod et al., 2024; Adams et al., 2024, grey circles).
@@ 2.3 UV luminosity functions
Because Henriques et al. (2015, 2020) tuned the L-Galaxies parameters to the low-redshift Universe ($z<3$), it is necessary to adapt them to match the redshift regime that we are interested in (see Vani et al., 2024, on discrepancies from observations at higher redshifts). In particular, for the LOFAR EoR KSP team, focussing on $8.510>z>$ 8.5 is crucial, as this is the regime in which the LOFAR telescope is most sensitive. Thus, we change the values of the parameters listed in Section 2.2.1 to match the observations of the UVLFs from HST legacy fields and JWST programs (Finkelstein et al., 2015; Bouwens et al., 2015, 2021; Harikane et al., 2022; Bouwens et al., 2023a; Bouwens et al., 2023b; Harikane et al., 2023; Leung et al., 2023; McLeod et al., 2024; Adams et al., 2024) at $z=10$ and 9. We have also varied other parameters, but their impact on the UVLFs is minimal (see M23, for a detailed analysis of the impact of different parameters). In future work, we will broaden the range of astrophysical parameters and analyse their effect on other observables.
First, we tune the $\alpha_{\rm SF}$, $\alpha_{\rm SF,burst}$, and $E_{\rm SN}$ values in order for the fiducial case to match the observed high-$z$ UVLFs. Additionally, we reduce the value of the fraction of AGN formed $(k_{\rm AGN})$ by a small amount, which improves the agreement at the brightest magnitudes ($M_{\rm 1600,AB}<-20$). Similarly, the supernova feedback-based heating efficiency parameters $\epsilon_{\rm reheat}$ and $V_{\rm reheat}$ are changed to improve the agreement for $M_{\rm 1600,AB}>-19$.
We consider two more cases. In the first one, the same L-Galaxies parameters of the fiducial model above are adopted for all $N$-body simulations, to investigate the impact of different cosmologies on the 21-cm signal independently from observations of galaxies at high redshifts. We refer to this model as “unconstrained”. This case allows us to have a clearer picture of the impact of changing cosmologies on the 21-cm signal, without astrophysical processes potentially cancelling out their effects on observables. In the second case, called “constrained”, the L-Galaxies parameters (specifically $\alpha_{\rm SF}$, $\alpha_{\rm SF,burst}$, and $E_{\rm SN}$) are changed for each cosmological model to match the observed UVLFs. This is potentially more interesting, as it allows for a joint constraint of astrophysical and cosmological parameters based on multi-frequency observations. The L-Galaxies parameters adopted are listed in Table 2.
@@
We show the UVLFs for the unconstrained (top) and constrained case (bottom) in Figure 2 for all cosmological models, along with observations at $z=10$ and 9. In the unconstrained case, we note that the fiducial and $h$ high models have similar UVLFs, while the $\sigma_{8}$ low model underpredicts the UVLF as compared to observations. On the other hand, the UVLF in the $\sigma_{8}$ high model is mildly overpredicted. These results are in agreement with the impact that these parameters are expected to have on matter clustering. In the constrained case, the boost in star formation due to higher $\alpha_{\rm SF}$ and $\alpha_{\rm SF,burst}$ values and a lower $E_{\rm SN}$ value, allows the $\sigma_{8}$ low model to match the bright end of the UVLF. However, due to lower matter clustering, it does not produce as many faint galaxies as the other cases, and sees a steep drop at $M_{\rm 1600,AB}>-18$. On the other hand, the high $E_{\rm SN}$ in the $\sigma_{8}$ high model strongly suppresses star formation and thus matches other cosmological models, as well as JWST and HST observations. Additionally, it produces more faint galaxies even at $M_{\rm 1600,AB}>-16$. Between the $h$ high and fiducial models, the difference is minimal at $z=10$ due to the mildly higher $E_{\rm SN}$ in the $h$ high model. However, at $z=9$ the impact of higher matter clustering in the $h$ high model shows up in the form of more galaxies at the low luminosity end.
@@
Lastly, we note that in the fiducial model the brightest galaxy at $z=14$ has $M_{\rm 1600,AB}=-20.36$, while at $z=13$ it has $M_{\rm 1600,AB}=-20.60$, which is in agreement with the recent spectroscopically confirmed galaxies at $z=14.32$ and $13.90$ (Carniani et al., 2024). For completeness, in Appendix A we also show the UVLFs in the range $512>z>$ 5 for the constrained case. We find that our models reasonably agree with UVLF observations across this redshift range.
@@ 2.4 Reionization history
Before discussing the reionization history in more detail, we note that the escape fraction of UV photons had been set to $f_{\rm esc}=12.5\%$ in order for the fiducial model to reionize completely (i.e. the average neutral fraction $\langle x_{\rm HI}\rangle$ becomes zero) by $z=5$. For the other models, though, full reionization is reached at different redshifts. In the future we plan to consider $f_{\rm esc}$ as a free parameter as well, although, as a reference, in Appendix C we show the reionization history of the fiducial model with $f_{\rm esc}=25\%$.
@@
For a qualitative view of the reionization history, in Figure 3 we show the A-series (i.e. with “fixed” initial conditions in the F&P pair of simulations) middle slices of the $\delta T_{\rm b}$ cubes for the four cosmological models in the constrained case at $z=12,10,8~{}\rm and~{}6$. We note that there are no major differences until $z=8$, when more and larger ionized regions appear in the $\sigma_{8}$ low model, while less and smaller ones are present in the $\sigma_{8}$ high model. This leads to the $\sigma_{8}$ low model being completely ionized by $z=6$, while the $\sigma_{8}$ high model still has large neutral regions. This is because, contrary to the fiducial model, in these cases the choice of the parameters controlling the astrophysical processes has an impact stronger than the one of $\sigma_{8}$ on the matter clustering. However, the differences between the astrophysical parameters of the $h$ high and the fiducial model are smaller, so that the differences in the size of the ionized regions from $z=10$ is mainly driven by matter clustering. Nevertheless, the choice of both astrophysical and cosmological parameters leads to significantly different speeds at which reionization happens across all four models. Indeed, despite the relatively similar maps observed at $z=8$, by $z=6$ the differences are much more evident. For completeness, in Appendix B we also show the B-series middle slices.
@@
In Figure 4, we present the redshift evolution of the average of the A and B-series volume-averaged neutral hydrogen fraction, $\langle x_{\rm HI}\rangle$, in the unconstrained (top) and constrained (bottom) cases, together with observational constraints. As in some models the redshift of reionization is not properly captured due to the coarse redshift resolution [^6] While we have a large number of snapshots as discussed in Section 2.1 for the sake of simplicity, to run Polar we adopt a uniform step size of $\Delta z=0.5$., we additionally provide a fit to these curves to estimate more accurately the end of reionization. We note that in the unconstrained case, all models reionize at different redshifts. Reionization is the fastest (at $z=6.4$) in the $\sigma_{8}$ high model, as structure formation happens earlier. The $\sigma_{8}$ low model is still only 45% reionized by $z=5$ because of the lack of sources, and shows the poorest agreement with observational constraints. In the constrained case, the differences are reduced, but still significant. In particular, we note that the $\sigma_{8}$ low model reionizes the fastest at $z=5.75$, while the $\sigma_{8}$ high model is only $\approx 65\%$ ionized by $z=5$. This reversal of reionization histories with respect to the unconstrained case is due to the impact of different astrophysical parameters. Specifically, the significantly higher $E_{\rm SN}$ in the $\sigma_{8}$ high model blows away gas and thus suppresses star formation, which consequently reduces the production of ionizing photons. On the other hand, the higher $\alpha_{\rm SF}$ and $\alpha_{\rm SF,burst}$ in the $\sigma_{8}$ low model along with the lower $E_{\rm SN}$ significantly boosts star formation.
@@
[…]

### v2

@@ 2.2.1 Semi-Analytic modeling of galaxies
We use L-Galaxies as described by Barrera et al. (2023) as a post-processing module of Gadget-4. This is an updated version of the publicly available L-Galaxies 2020 that was used by M23. L-Galaxies implements most major physical processes of gas cooling, star formation, galaxy mergers, supernovae feedback, black hole growth, AGN feedback, and dust attenuation. While Barrera et al. (2023) largely builds on Henriques et al. (2015), we also consider the parameters used by Henriques et al. (2020) before adapting them for our purposes. In particular, we focus on those that scale the star formation efficiency ($\alpha_{\rm SF}$), star formation efficiency during galaxy mergers ($\alpha_{\rm SF,burst}$), AGN accretion rate ($k_{\rm AGN}$), reheating of cold gas by star formation ($\epsilon_{\rm reheat}$, $V_{\rm reheat}$), and the energy released by each supernova ($E_{\rm SN}$). L-Galaxies assumes that stars form from the cold gas within the disk of each galaxy. Thus, these parameters control the amount of cold gas available in the disk and the efficiency of converting it into stars. For more details, refer to the supplementary material of Henriques et al. (2015).
@@ 2.2.2 Radiative transfer and the 21-cm signal
More specifically, Grizzly requires as input the gridded matter density field and dark matter halo masses from the $N$-body simulation, as well as the corresponding galactic stellar masses and stellar ages obtained with L-Galaxies. Our reference simulation has a grid of 2563 cells, resulting in a cell size of $\approx 600~h^{-1}~\rm ckpc$ [^7] The grid size is chosen after checking the convergence of the volume-averaged neutral hydrogen fraction and the 21-cm power spectrum in the fiducial model for a range of grid sizes (from 643 to 5123 cells in steps of 23) across redshifts $12>z>5$. We note that the values are within 5%, and thus opt for 2563 to ensure we do not miss out on the impact of low-density regions in any redshift regime.. Next, we assume that the gas density scales by a factor of $\Omega_{\rm b}/\Omega_{\rm dm}$ with the dark matter density. Lastly, as done in M23, we use the Binary Population and Spectral Synthesis (BPASS; Stanway & Eldridge, 2018) code to model the Spectral Energy Distributions (SEDs) of stellar sources. In future work, we will also explore the impact of other source types such as X-ray binaries, shock-heated interstellar medium, and accreting black holes (as done in Eide et al., 2018, 2020; Ma et al., 2021).
@@
where $x_{\rm HI}$ is the fraction of neutral hydrogen, $\delta_{\rm B}$ is the fractional overdensity of baryons, $T_{\rm S}$ is the hydrogen spin temperature, $T_{\rm CMB}$ is the temperature of the CMB photons at redshift $z$, $h$ is the dimensionless Hubble constant, and $\Omega_{\rm b}$ and $\Omega_{\rm m}$ are the baryonic matter density and total matter density parameters respectively. As done by M23, we assume $T_{\rm S}\gg T_{\rm CMB}$ which is valid when the IGM has been sufficiently heated by X-ray sources and expected to be the case in the range of redshift of interest here. Additionally, we also ignore the impact of redshift space distortions. Thus Equation 1 is reduced to
@@
We define the power spectrum of $\delta T_{\rm b}$ as (Mertens et al., 2020; Metha & Berger, 2024):
@@
where $\delta_{D}$ is the Dirac delta function, $\delta T_{\rm b}(\mathbf{k})$ is the differential brightness temperature in Fourier space, and $\langle$$\rangle$ is the ensemble average. In the following, we will report our results in terms of the normalized form of the power spectrum, given by (see Peacock, 1999, for details):
@@
Table 2: List of astrophysical parameters in L-Galaxies (see Henriques et al., 2015, 2020, for more details) that can be tuned to control galaxy formation and evolution. We show the names of the parameters in column 1, and Henriques et al. (2015) and Henriques et al. (2020) values tuned to $z<3$ observations in columns 2 and 3. Our cosmological models (fiducial, $h$ high, $\sigma_{8}$ low and $\sigma_{8}$ high) tuned to UVLF observations at $z=10$ and $9$ are shown in columns 4, 5, 6 and 7 respectively. We highlight the parameters that differ from Henriques et al. (2015, 2020) values in bold and fix the values of the unchanged parameters to the latter. For the unconstrained case (see text for more details), we use the fiducial model parameters for all our NN-body simulations.
@@
$Z_{\rm yield}$ 0.046 0.030 0.030 0.030 0.030 0.030
Figure 2: A and B-series averaged UVLFs for the unconstrained (top) and constrained (bottom) cases at $z=10$ and 9 for the fiducial (cyan), $h$ high (red), $\sigma_{8}$ low (purple) and $\sigma_{8}$ high (green). The averaging of the F&P pairs allows to compensate for stochasticity in underfilled bins of the UVLFs, by suppressing the impact of cosmic variance. We also show the dust attenuated (solid) and unattenuated (dashed) UVLFs, and JWST and HST observations (Finkelstein et al., 2015; Bouwens et al., 2015, 2021; Harikane et al., 2022; Bouwens et al., 2023a; Bouwens et al., 2023b; Harikane et al., 2023; Leung et al., 2023; McLeod et al., 2024; Adams et al., 2024, grey circles).
@@ 2.3 UV luminosity functions
Because Henriques et al. (2015, 2020) tuned the L-Galaxies parameters to the low-redshift Universe ($z<3$), it is necessary to adapt them to match the redshift regime that we are interested in (see Vani et al., 2024, on discrepancies from observations at higher redshifts). In particular, for the LOFAR EoR KSP team, focussing on $10>z>8.5$ is crucial, as this is the regime in which the LOFAR telescope is most sensitive. Thus, we change the values of the parameters listed in Section 2.2.1 to match the observations of the UVLFs from HST legacy fields and JWST programs (Finkelstein et al., 2015; Bouwens et al., 2015, 2021; Harikane et al., 2022; Bouwens et al., 2023a; Bouwens et al., 2023b; Harikane et al., 2023; Leung et al., 2023; McLeod et al., 2024; Adams et al., 2024) at $z=10$ and 9 by employing the least square minimisation [^8] In the least square minimisation the square of the residuals between UVLF observations and models is summed for both redshifts over a range of different values for the parameters. The values that lead to the minimum sum of the squares of the residuals are accepted.. We iteratively vary parameters one at a time, with the range going from 1/100th of the lower value to 100 times the higher value between the fitted values obtained by Henriques et al. (2015) and Henriques et al. (2020) for each parameter. While typically the average energy released by each supernova is expected to be $\approx 10^{51}$ ergs (see Jones et al., 1998), we find that a mild reduction to $0.8\times 10^{51}$ ergs provides a better fit, while still being consistent with observations (Rubin et al., 2016; Kozyreva et al., 2025). Further, the resulting values for $\alpha_{\rm SF}$ and $\alpha_{\rm SF,burst}$ suggest higher star formation efficiency, in agreement with past observations (Stefanon et al., 2021), abundance matching estimates (Tacchella et al., 2018) and models (Sun & Furlanetto, 2016; Sipple & Lidz, 2024). The mild reduction of $k_{\rm AGN}$ with respect to the values in Henriques et al. (2020) leads to lower gas accretion by AGNs, allowing for more star formation and thus more galaxies at the brightest magnitudes ($M_{\rm 1600,AB}<-20$), which is necessary to match recent observations. Similarly, the small differences in the supernova feedback-based heating efficiency parameters $\epsilon_{\rm reheat}$ and $V_{\rm reheat}$ are required to improve the agreement for $M_{\rm 1600,AB}>-19$. Indeed, as supernovae can effectively disrupt smaller galaxies, reducing the impact of their heating efficiency allows for more cold gas to form, and thus more star formation, further increasing the UVLF at the low brightness end and providing a closer match with observations.
We have also varied other parameters affecting star formation ($\beta_{\rm SF,burst}$), the heating of cold gas by stars and supernova feedback (such as $\beta_{\rm reheat}$, $\eta_{\rm eject}$, $V_{\rm eject}$, $\beta_{\rm eject}$, $\gamma_{\rm reinc}$), black hole properties ($f_{\rm BH}$, $V_{\rm BH}$), and galaxy mergers ($M_{\rm r.p.}$, $\alpha_{\rm dyn.fric.}$), but their impact on the UVLFs is minimal (see M23, for a detailed analysis of the impact of different parameters). For such parameters, we fix their value to that obtained by Henriques et al. (2020). In future work, we will broaden the range of astrophysical parameters and analyse their effect on other observables. Additionally, we also plan to employ more rigorous MCMC-based techniques for fitting parameters (as shown in Ma et al., 2025, submitted).
So far, we only considered the fiducial model. While the values of the parameters obtained may be sensitive to the method used to derive them (in this case, the least squares minimization), the qualitative difference in their values for the different cosmological models introduced in Section 2.1 can still be explored. Thus, we consider two cases. In the first one, the same L-Galaxies parameters of the fiducial model above are adopted for all $N$-body simulations to investigate the impact of different cosmologies on the 21-cm signal independently from observations of galaxies at high redshifts. We refer to this model as “unconstrained”. This case enables us to gain a clearer understanding of the impact of changing cosmologies on the 21-cm signal, without astrophysical processes potentially cancelling out their effects on observables. However, not changing the astrophysical parameters for the different cosmological models may produce UVLFs that are inconsistent with observations. Thus, we also consider another scenario, called the “constrained” case. Here, the L-Galaxies parameters (specifically $\alpha_{\rm SF}$, $\alpha_{\rm SF,burst}$, and $E_{\rm SN}$) are changed for each cosmological model to match the observed UVLFs. As the average energy released per supernova is the most sensitive parameter in scaling the UVLF while maintaining its overall shape, we first focus on tuning it. For the $h$ high model, we find that fixing it at $10^{51}$ ergs and using the fiducial model’s values for $\alpha_{\rm SF}$ and $\alpha_{\rm SF,burst}$ is enough to get the best fit via the least squares minimisation described above. The increase in $E_{\rm SN}$ as compared to the fiducial model is required to reduce the amount of cold gas in the galaxy available for star formation, balancing the increased clustering of matter due to the higher Hubble constant. Similarly, due to even stronger matter clustering due to a larger $\sigma_{8}$ in the $\sigma_{8}$ high model, we note that a value of $2\times 10^{51}$ ergs for $E_{\rm SN}$ provides the best fit. Because of this, we expect that just reducing $E_{\rm SN}$ for the $\sigma_{8}$ low model to improve star formation efficiency is enough despite the lower matter clustering. However, we note that in addition to reducing the average supernova energy to $0.15\times 10^{51}$ ergs, it is also necessary to increase the parameters regulating the star formation efficiency, $\alpha_{\rm SF}$, and the burstiness of star formation efficiency during mergers, $\alpha_{\rm SF,burst}$. Further raising the values beyond 0.50 for $\alpha_{\rm SF}$ and 0.90 for $\alpha_{\rm SF,burst}$ does not induce any appreciable difference in the sum of the squares of the residuals.
These differences in the astrophysical parameters for various cosmologies are potentially more interesting, as they allow for a joint constraint of astrophysical and cosmological parameters based on multi-frequency observations. The list of all the L-Galaxies parameters adopted for the four models is provided in Table 2.
@@
We show the UVLFs for the unconstrained (top) and constrained (bottom) case in Figure 2 for all cosmological models, along with observations at $z=10$ and 9. In the unconstrained case, we note that the fiducial and $h$ high models have similar UVLFs, while the $\sigma_{8}$ low model underpredicts the UVLF as compared to observations. On the other hand, the UVLF in the $\sigma_{8}$ high model is mildly overpredicted. These results are in agreement with the impact that the parameters are expected to have on matter clustering, as discussed above. In the constrained case, the boost in star formation due to higher $\alpha_{\rm SF}$ and $\alpha_{\rm SF,burst}$ values and a lower $E_{\rm SN}$ value, allows the $\sigma_{8}$ low model to match the bright end of the UVLF. However, due to lower matter clustering, it does not produce as many faint galaxies as the other cases, and sees a steep drop at $M_{\rm 1600,AB}>-18$. On the other hand, the high $E_{\rm SN}$ in the $\sigma_{8}$ high model strongly suppresses star formation and thus matches other cosmological models, as well as JWST and HST observations. Additionally, it produces more faint galaxies even at $M_{\rm 1600,AB}>-16$. Between the $h$ high and fiducial models, the difference is minimal at $z=10$ due to the mildly higher $E_{\rm SN}$ in the $h$ high model. However, at $z=9$ the impact of higher matter clustering in the $h$ high model shows up in the form of more galaxies at the low luminosity end.
@@
Lastly, we note that in the fiducial model the brightest galaxy at $z=14$ has $M_{\rm 1600,AB}=-20.36$, while at $z=13$ it has $M_{\rm 1600,AB}=-20.60$, which is in agreement with the recent spectroscopically confirmed galaxies at $z=14.32$ and $13.90$ (Carniani et al., 2024). For completeness, in Appendix A we also show the UVLFs in the range $12>z>5$ for the constrained case. We find that our models reasonably agree with UVLF observations across this redshift range.
@@ 2.4 Reionization history
Before discussing the reionization history in more detail, we note that the escape fraction of UV photons had been set to $f_{\rm esc}=12.5\%$ in order for the fiducial model to reionize completely (i.e. the average neutral fraction $\langle x_{\rm HI}\rangle$ becomes zero) by $z=5$. For the other models, though, full reionization is reached at different redshifts. Further, we note that this is a simplified assumption, as $f_{\rm esc}$ is expected to vary with galaxy and halo properties (as shown by Kostyuk et al., 2023; Ma et al., 2020) and maybe with redshift (while e.g. Ma et al. 2020 and Kostyuk et al. 2025 find redshift variation, Gorce et al. 2018 favours $f_{\rm esc}$ to be constant with redshift), and thus can be strongly affected by modelling choices and approximations. Nevertheless, our choice aligns with recent estimates from observations (Mascia et al., 2023; Bosman & Davies, 2024; Jaskot et al., 2024, which report median values of 10-12%). In the future, we plan to consider $f_{\rm esc}$ as a free parameter as well, and, as a reference, in Appendix B we show the reionization history of the fiducial model with $f_{\rm esc}=25\%$. This is of particular interest, as observations like Jaskot et al. (2024) have reported higher escape fractions ($f_{\rm esc}>20\%$) for $\geq 33\%$ of observed galaxies.
@@
For a qualitative view of the reionization history, in Figure 3 we show the A-series (i.e. with “fixed” initial conditions in the F&P pair of simulations) middle slices of the $\delta T_{\rm b}$ cubes for the four cosmological models in the constrained case at $z=12,10,8~\rm and~6$. For completeness, in Appendix C.1 we also show the corresponding B-series slices, and in Appendix C.2 both the A and B-series middle slices for the unconstrained case. We note that there are no major differences until $z=8$, when more and larger ionized regions appear in the $\sigma_{8}$ low model, while fewer and smaller ones are present in the $\sigma_{8}$ high model. This leads to the $\sigma_{8}$ low model being almost completely ionized by $z=6$ (note that while in Figure 3, the shown slice at $z=6$ happens to be fully ionized, there are still some neutral regions in the full simulation box. The same is shown below by computing the volume-averaged neutral hydrogen fraction), while the $\sigma_{8}$ high model still has large neutral regions. This is because in these cases the choice of the parameters controlling the astrophysical processes at high redshift is balanced by the impact on the matter clustering due to the choice of $\sigma_{8}$, but at lower redshift, once enough halos have formed, the astrophysical processes dominate. In particular, a higher (lower) average energy per supernova in the $\sigma_{8}$ high (low) case significantly impedes (bolsters) star formation as too much cold gas is blown away (collapsed). However, the differences between the astrophysical parameters of the $h$ high and the fiducial model are smaller, so the differences in the size of the ionized regions from $z=10$ are mainly driven by matter clustering. Thus, overall, the choice of both astrophysical and cosmological parameters leads to significantly different speeds at which reionization happens across all four models. Indeed, despite the relatively similar maps observed at $z=8$, by $z=6$ the differences are much more evident.
@@
In Figure 4, we present the redshift evolution of the average of the A and B-series volume-averaged neutral hydrogen fraction, $\langle x_{\rm HI}\rangle$, in the unconstrained (top) and constrained (bottom) cases, together with observational constraints. As in some models the redshift of reionization is not properly captured due to the coarse redshift resolution [^9] While we have a large number of snapshots as discussed in Section 2.1 for the sake of simplicity, to run Polar we adopt a uniform step size of $\Delta z=0.5$., we additionally provide a fit to these curves (shown as a dotted line) to estimate more accurately the end of reionization. We note that in the unconstrained case, all models reionize at different redshifts. Reionization is the fastest (at $z=6.4$) in the $\sigma_{8}$ high model, as structure formation happens earlier. The $\sigma_{8}$ low model is still only 45% reionized by $z=5$ because of the lack of sources, and shows the poorest agreement with observational constraints. In the constrained case, the differences are reduced, but still significant. In particular, we note that the $\sigma_{8}$ low model reionizes the fastest at $z=5.75$, while the $\sigma_{8}$ high model is only $\approx 65\%$ ionized by $z=5$. This reversal of reionization histories with respect to the unconstrained case is due to the impact of different astrophysical parameters. Specifically, the significantly higher $E_{\rm SN}$ in the $\sigma_{8}$ high model blows away gas and thus suppresses star formation, which consequently reduces the production of ionizing photons. On the other hand, the higher $\alpha_{\rm SF}$ and $\alpha_{\rm SF,burst}$ in the $\sigma_{8}$ low model along with the lower $E_{\rm SN}$ significantly boosts star formation.
@@
Finally, the four models produce a Thomson scattering optical depth in the range (0.041-0.067) and (0.048- 0.060) for the unconstrained and constrained case respectively, assuming an Helium (He) II fraction equal to the HII fraction, and an instantaneous HeII reionization at z= 3z\,=\,3. These numbers are in agreement with Planck observations (Planck Collaboration et al., 2020; de Belsunce et al., 2021; Giarè et al., 2024).